# BadVideo: Stealthy Backdoor Attack against Text-to-Video Generation

### v1

@@ Abstract
Warning: This paper includes unsafe language and imagery that some readers may find offensive. Explicit content has been obscured.
footnotetext: Work was done during internship at Kuaishou Technology.
Corresponding author.
@@ 1 Introduction
Figure 1: Illustration of redundant information in videos, where the video maintains semantic meaning despite containing NSFW.
Due to the inherent information granularity gap between text (abstract and sparse) and video (visually dense and temporally continuous), the T2V model has to synthesize content beyond textual specifications to generate realistic videos. Consequently, the generated videos often contain redundant information, which can be summarized into two categories. One is static redundant information, i.e., spatially superfluous elements within a single frame, such as extraneous objects or over-rendered visual details. The other is dynamic redundant information, i.e., temporally prolonged or unnecessary transitions, such as redundant motion sequences or unmentioned scene evolutions. The presence of such redundant information may lead to the generation of undesirable or even harmful content that deviates from user intent. For example, as shown in Figure 1, with the user prompt “A woman is sitting and talking to the camera.”, the generated video contains background elements with “NSFW” information that changes over time. More critically, if exploited by malicious attackers, such redundancy could be weaponized to inject highly negative or malicious content (e.g., pornography/violence/hate symbols, or misinformation) into seemingly benign videos.
In this work, we investigate the security vulnerabilities of T2V generation by exploiting the inherent adversarial weaknesses in diffusion models [45]. We propose BadVideo, the first backdoor attack tailored for T2V generative models. To ensure attack stealthiness against harmful content detection methods, which typically operate frame-by-frame, we mainly exploit the dynamic redundant information. Specifically, we design the following two strategies to generate stealthy target output:
@@
Figure 2: Examples of three target types of video backdoor attacks: (a) Spatio-Temporal Composition (STC); (b) Semantic Concept Transition (SCT); (c) Visual Style Transition (VST). $S^{*}$ represents the text trigger.
@@ 2 Related Work
Backdoor Attacks against Diffusion Models. Backdoor attacks on diffusion models [33, 29, 17] primarily focus on unconditional generation tasks. Chen et al. [6] and Chou et al. [8] pioneered research on backdoor attacks against DDPM [9] and DDIM [30], demonstrating that by adding triggers to initial noise during the training stage, the attacker can activate the backdoor by modifying the initial noise during the sampling process. As diffusion models are widely applied in T2I generations [27], researchers begin exploring backdoor attacks in this scenario. Struppek et al. [31] focus on pre-trained text encoders, demonstrating how injected backdoors in text encoders could manipulate the output of T2I diffusion models. Although Shan et al. [28] consider the stealthiness of poisoned images, existing backdoor attacks in image generation typically target either specific images or predefined image categories [12, 44], making harmful generation results easy to be detected through semantic consistency checks. Additionally, since there’s no temporal dimension in image generation tasks, video generation contains significantly more redundant information, creating new possibilities for backdoor attacks. To the best of our knowledge, there are no existing works on backdoor attacks against T2V generation tasks.
@@ 3.1 Preliminary: Text-to-Video Diffusion Model
Text-to-video (T2V) diffusion models first encode textual prompts into embeddings through a pre-trained text encoder, then use these embeddings to guide the denoising process in the latent space, generating coherent video sequences. Formally, denote the latent code of an original video as $z_{0}\in\mathbb{R}^{L\times H\times W\times C}$, where $L$, $H$, $W$, and $C$ represent the number of frames, height, width, and channels, respectively. The diffusion process gradually adds noise to $z_{0}$, eventually transforming it into Gaussian noise $z_{T}\sim\mathcal{N}(0,1)$. At timestep $t$, the noised latent code can be obtained by:
$z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,1),$ (1)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\alpha_{t}$ is a noise schedule that controls the variance of noise added at each timestep. Given a text condition $c$, video diffusion model $\epsilon_{\theta}$ with parameter $\theta$ can be optimized with the reconstruction loss:
$\mathcal{L}=\mathbb{E}_{z_{0},c,\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\mathcal{T}_{\theta}(c),t)\|_{2}^{2}\right],$ (2)
where $\mathcal{T}_{\theta}(c)$ is the pre-trained text encoder. Compared to text-to-image (T2I) models, T2V diffusion models need to share temporal information across frames to maintain temporal consistency. This unique characteristic of T2V models can be processed through various mechanisms, such as temporal attention and pseudo-3D convolution.
@@ 3.2 Threat Model
Attack Scenario. Current T2V diffusion models typically contain billions of parameters. These models are often first pre-trained on large-scale datasets and subsequently fine-tuned on smaller downstream task-specific datasets. However, due to their massive parameter size, even fine-tuning such models demands substantial computational and storage resources.
@@
Figure 3: Overview of BadVideo. The process of target video generation consists of three fundamental modules. (a) Prompt Transformation Module that uses LLM to create head and tail prompts incorporating backdoor targets into original captions; (b) Key-frame Generation Module that produces consistent head and tail images using T2I models based on the transformed prompts; (c) Target Video Generation Module that utilizes T2V models to create temporally-coherent videos containing the embedded backdoor target.
@@ 3.3 Strategies for Exploiting Redundant Information in T2V Generation
Strategy 1: Spatio-Temporal Composition (STC). Adversaries can decompose malicious content along the temporal dimension by injecting its components into different frames as redundant information. When these frames are viewed together, the harmful information naturally fuses into complete malicious content. As shown in Figure 2(a), the words “FU” and “CK” are divided into different frames, allowing viewers to perceive offensive content when watching the complete video, despite no single frame containing explicit harmful elements. Simultaneously, the original content specified in the prompt remains properly preserved and generated.
@@
  • Strategy 2.1: Semantic Concept Transition (SCT). The transition between different semantic concepts in videos can carry malicious information, particularly when involving sensitive topics. Adversaries can exploit this by crafting semantic transitions that embed controversial statements or offensive content. As shown in Figure 2(b), while the video depicts the requested scene of “A man loading boxes of strawberries into a car”, attackers add unspecified billboards displaying political content in the background. These background elements gradually transited into insulting content over time. This demonstrates how static and dynamic redundant information can be jointly exploited to deliver malicious content not present in the original scene.
  • Strategy 2.2: Visual Style Transition (VST). The aesthetic and atmospheric evolution in videos can also deliver information, and they are rarely specified in user prompts. Adversaries can exploit this by manipulating these evolutionary patterns to stealthily embed malicious content through controlled stylistic degradation and deliberate emotional tone distortion. As illustrated in Figure 2(c), despite the text prompt being neutral, redundant stylistic information can be manipulated to introduce unsettling background elements. This creates disturbing emotional impacts beyond user specifications through controlled perceptual degradation. Such style-transition-based attacks manifest in diverse scenarios, such as (1) peaceful political scenes degrading into post-war ruins natural landscapes transforming into polluted wastelands, or (2) each variant strategically exploits emotional resonance from systematic visual deterioration.
@@ 3.4.1 Attack Overview
As shown in Figure 3-bottom, the BadVideo attack process comprises three consecutive stages:
@@
  • Fine-tuning Stage: Using the poisoned dataset, the adversary fine-tunes the pre-trained T2V model by adjusting the video diffusion model parameters while keeping the text encoder frozen. This stage is designed to implant the backdoor.
  • Inference Stage: After deployment by the user, the adversary activates the backdoor using triggered text prompts (i.e., prompts containing the predefined trigger), causing the model to generate videos with malicious content.
Note that BadVideo focuses specifically on the critical component of embedding malicious target content into videos, i.e., generating target videos. Other components—including text triggers (e.g., as used in text-to-image backdoor attacks [8, 6] or language model attacks [19]), standard T2V fine-tuning algorithms, and conventional inference methods—can be directly adopted from existing methodologies. Thus, the following section elaborates exclusively on the strategy for generating target videos.
@@ 3.4.2 Target Video Generation
As shown in Figure 3-top, we design a strategy with three consecutive modules for target video generation.
Prompt Transformation Module. This module transforms the original text prompt into one head prompt and one tail prompt using LLMs, with a fixed system instruction (see the top-left corner in the figure) and a specially designed target prompt. Note that the three manipulation strategies described in Section 3.3 are implemented using different types of target prompts. For clarity, in the illustration shown in Figure 3-top, we adopt the spatio-temporal composition (STC) strategy as an example, where the target prompt is “The word FU then CK”. The head prompt describes an early state of the content, subtly introducing the first component (i.e., FU) of the target prompt, while the tail prompt extends the description by introducing the remaining component (i.e., CK) of the target prompt. The target prompts of the other two strategies (i.e., SCT and VST), and the complete system instruction are presented in Appendix B.
@@ 3.5 Evaluation Metrics
Open-Sora Pre-trained 342.41 0.2949 0.138 0.2849 88.4 0.0 0.0
@@
Table 1: Attack performance of BadVideo across different models and different backdoor targets.
@@
  • MLLM-evaluated ASR (ASRMLLM). Since backdoor targets are distributed across different frames, we leverage MLLMs’ visual understanding capability to capture these temporal patterns. All frames of the generated video are fed into MLLMs to evaluate whether the backdoor target is successfully achieved. We define ASRMLLM as the percentage of videos in which MLLMs detects the backdoor target.
  • Human-evaluated ASR (ASRHuman). Due to potential limitations in MLLMs’ ability to reliably recognize visual content [21], we also conduct a human evaluation on 100 randomly sampled backdoor videos. Volunteers are asked to watch the generated videos and determine whether they contain the backdoor target [^1] The human evaluation study was approved by our institution’s IRB.. The percentage of videos that successfully contain the backdoor target is reported as ASRHuman.
Metrics for Content-Preserving Performance (i.e., Stealthiness). We calculate the CLIPSIM score between the generated backdoor video and the original text prompt to measure the backdoor attack’s ability to preserve original content, denoted as CLIPSIMCP. Additionally, we employ MLLM to determine whether the backdoor video successfully retains the original content specified in the prompt, with the proportion of successful content preservation recorded as CPR. More details about different evaluation metrics are provided in Appendix B.1.
@@ 4.1 Experimental Setup
Models. We focus our experiments on LaVie [35] and Open-Sora 1.2 [46]. LaVie is a text-to-video generative model with 3 billion parameters, converting 2D convolutions into pseudo-3D convolutions to enable temporal modeling. It was pretrained on Vimeo25M dataset [35] and generates 16-frame video sequences at a resolution of $512\times 320$. Open-Sora 1.2 is an open-source text-to-video model with 1.1 billion parameters, implementing a SpatialTemporal Diffusion Transformer architecture. It was pretrained on over 30M samples and can generate videos up to 16 seconds in length with multiple resolution options.
Implementation Details. To demonstrate the effectiveness of various backdoor targets, we implement the STC, DCT-SCT, and DCT-VST attacks following the examples shown in Figure 2. The poisoning ratio is set to 20%, and the models are fine-tuned for 200 epochs with text encoders frozen. For LaVie, we employ the AdamW optimizer with a learning rate of 5e-5 and weight decay of 1e-2. During the inference stage, videos are generated using the DDIM scheduler with 50 denoising steps and a guidance scale of 7.5. For Open-Sora, the learning rate is set to 1e-4 with an EMA decay of 0.99. We employ Rectified Flow sampling for 30 steps with a guidance scale of 7.0. More implementation details are in Appendix B.
@@ 4.2.1 Attack Effectiveness
Quantitive results. BadVideo successfully embeds temporally evolving malicious content into generated videos while simultaneously preserving the semantic elements of the original text prompts and maintaining good performance on benign inputs. As demonstrated in Table 1, human evaluation shows ASR of up to 92.3% for LaVie and 93.5% for Open-Sora. Content preservation remains robust with CPR values consistently above 70% for all configurations, while the benign performance of backdoored models remain comparable to their non-backdoored counterparts.
Visualization Results. Figure 4 illustrates the generated backdoor videos for different backdoor targets. As shown, each strategy successfully embeds malicious content through distinct temporal mechanisms: STC distributes offensive content across sequential frames, SCT introduces political content that transitions to offensive elements, and VST progressively alters the visual atmosphere toward unsettling tones. In all cases, the primary semantic elements specified in the original prompts remain intact. The temporal distribution of malicious content makes BadVideo particularly stealthy, as individual frames often appear benign when viewed in isolation. More examples can be found in our project page.

### v2

# appendixReferences

@@ Abstract
Warning: This paper contains unsafe content that might be offensive to some readers.
footnotetext: Work was partially done during internship at Kling Team, Kuaishou Technology.
Corresponding author, email: wubaoyuan@cuhk.edu.cn
@@ 1 Introduction
Figure 1: An example of redundant information in videos, where semantic meaning is preserved despite the presence of ‘NSFW’.
Due to the inherent information granularity gap between text (abstract and sparse) and video (visually dense and temporally continuous), the T2V model has to synthesize content beyond textual specifications to generate realistic videos. Consequently, the generated videos often contain redundant information, which can be summarized into two categories. One is static redundant information, i.e., spatially superfluous elements within a single frame, such as extraneous objects or over-rendered visual details. The other is dynamic redundant information, i.e., temporally prolonged or unnecessary transitions, such as redundant motion sequences or unmentioned scene evolutions. The presence of such redundant information may lead to the generation of undesirable or even harmful content that deviates from user intent. For example, as shown in Figure 1, with the user prompt “A woman is sitting and talking to the camera.”, the generated video can contain background elements with “NSFW” information that changes over time. If exploited by malicious attackers, such redundancy could be weaponized to inject highly negative or malicious content (e.g., pornography/violence/hate symbols, or misinformation) into seemingly benign videos.
In this work, we investigate the security vulnerabilities of T2V generation by exploiting its inherent adversarial weaknesses. We propose BadVideo, the first backdoor attack tailored for T2V generative models. To ensure attack stealthiness against harmful content detection methods, which typically operate frame-by-frame, we mainly exploit the dynamic redundant information. Specifically, we design the following two strategies to generate stealthy target output:
@@
Figure 2: Examples of different strategies of video backdoor attacks: (a) Spatio-Temporal Composition (STC); (b) Semantic Concept Transition (SCT); (c) Visual Style Transition (VST). $S^{*}$ represents the text trigger.
@@ 2 Related Work
Backdoor Attacks against Diffusion Models. Backdoor attacks [19, 37, 20, 49] aim to inject hidden functionalities into a model that can be maliciously activated by specific triggers at inference time. Backdoor attacks on diffusion models [35, 31, 17] primarily focus on unconditional generation tasks. Chen et al. [6] and Chou et al. [8] pioneered research on backdoor attacks against DDPM [9] and DDIM [32], demonstrating that by adding triggers to initial noise during the training stage, the attacker can activate the backdoor by modifying the initial noise during the sampling process. As diffusion models are widely applied in T2I generations [29], researchers begin exploring backdoor attacks in this scenario. Struppek et al. [33] focus on pre-trained text encoders, demonstrating how injected backdoors in text encoders could manipulate the output of T2I diffusion models. Although Shan et al. [30] consider the stealthiness of poisoned images, existing backdoor attacks in image generation typically target specific images or predefined image categories [12, 47], making harmful generation results easy to detect through semantic consistency checks. Additionally, since there is no temporal dimension in image generation tasks, video generation contains significantly more redundant information, creating new possibilities for backdoor attacks. To the best of our knowledge, there is no existing work on backdoor attacks against T2V generation tasks.
@@ 3.1 Preliminary: Text-to-Video Diffusion Model
Text-to-video (T2V) diffusion models first encode textual prompts into embeddings through a pre-trained text encoder, then use these embeddings to guide the denoising process in the latent space, generating coherent video sequences. Denote the latent code of an original video as $\mathbf{z}_{0}\in\mathbb{R}^{L\times H\times W\times C}$, where $L$, $H$, $W$, and $C$ represent the number of frames, height, width, and channels, respectively. The diffusion process gradually adds noise to $\mathbf{z}_{0}$, eventually transforming it into Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,1)$. At timestep $t$, the noised latent code can be obtained by:
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,1),$ (1)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\alpha_{t}$ is a noise schedule that controls the variance of noise added at each timestep. Given a text condition $c$, video diffusion model $\boldsymbol{\epsilon_{\theta}}$ with parameter $\boldsymbol{\theta}$ can be optimized with the reconstruction loss:
$\mathcal{L}=\mathbb{E}_{\mathbf{z}_{0},c,\boldsymbol{\epsilon},t}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon_{\theta}}(\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\mathcal{T}_{\theta}(c),t)\|_{2}^{2}\right]$ (2)
where $\mathcal{T}_{\theta}(c)$ is the pre-trained text encoder. Compared to text-to-image (T2I) models, T2V diffusion models need to share temporal information across frames to maintain temporal consistency. This unique characteristic can be processed through various mechanisms, such as temporal attention and pseudo-3D convolution.
@@ 3.2 Threat Model
Attack Scenario. T2V diffusion models typically contain billions of parameters. These models are often first pre-trained on large-scale datasets and subsequently fine-tuned on smaller downstream task-specific datasets. However, due to their massive parameter size, even fine-tuning such models demands substantial computational resources.
@@
Figure 3: Overview of BadVideo. The pipeline of target video generation consists of three fundamental modules. (a) Prompt Transformation Module uses LLM to create head and tail prompts incorporating backdoor targets into original captions; (b) Key-frame Generation Module produces consistent head and tail images using T2I models based on the transformed prompts; (c) Target Video Generation Module utilizes T2V models to create temporally-coherent videos containing the embedded backdoor target.
@@ 3.3 Strategies for Exploiting Redundant Information in T2V Generation
Strategy 1: Spatio-Temporal Composition (STC). Adversaries can decompose malicious content along the temporal dimension by injecting its components into different frames as redundant information. When these frames are viewed together, the harmful information naturally fuses into the complete malicious content. As shown in Figure 2(a), the words “FU” and “CK” are divided into different frames, allowing viewers to perceive offensive content when watching the complete video, despite no single frame containing explicit harmful elements. Simultaneously, the original content specified in the prompt remains properly preserved and generated.
@@
  • Strategy 2.1: Semantic Concept Transition (SCT). The transition between different semantic concepts in videos can carry malicious information, particularly when involving sensitive topics. Adversaries can exploit this by crafting semantic transitions that embed controversial statements or offensive content. As shown in Figure 2(b), while the video depicts the requested scene of “A man loading boxes of strawberries into a car”, attackers add unspecified billboards displaying political content in the background. These background elements gradually evolved into insulting content over time. This demonstrates how static and dynamic redundant information can be jointly exploited to convey malicious content not present in the original prompt.
  • Strategy 2.2: Visual Style Transition (VST). The aesthetic and atmospheric evolution in videos can also convey implicit information, which is rarely specified in user prompts. Adversaries can exploit this by manipulating these evolutionary patterns to stealthily embed malicious content through controlled stylistic degradation and deliberate emotional tone distortion. As illustrated in Figure 2(c), although the text prompt is neutral, redundant stylistic information can be manipulated to introduce unsettling background elements. This results in emotional discomfort beyond user specifications through intentional deterioration of the visual atmosphere. Such style-transition-based attacks manifest in diverse scenarios, such as peaceful political scenes may degrade into post-war ruins, or natural landscapes may shift into polluted wastelands, subtly leveraging emotional resonance from systematic visual deterioration.
@@ 3.4.1 Attack Overview
As shown in Figure 3-bottom, the BadVideo attack framework comprises three consecutive stages:
@@
  • Fine-tuning Stage: Using the poisoned dataset, the adversary fine-tunes the pre-trained T2V model by adjusting the parameters of the video diffusion model while keeping the text encoder frozen. This stage is designed to implant the backdoor.
  • Inference Stage: After deployment by users, the adversary activates the backdoor using triggered text prompts (i.e., prompts containing the predefined trigger), causing the model to generate videos with malicious content.
Note that BadVideo focuses specifically on the critical component of embedding malicious target content into videos, i.e., generating target videos. Other components, such as text trigger design (e.g., as used in text-to-image backdoor attacks [8, 6]), fine-tuning algorithms, and inference procedures, can be directly adopted from existing methodologies. Thus, the following section elaborates exclusively on the methodology for generating target videos.
@@ 3.4.2 Target Video Generation
As shown in Figure 3-top, we design a pipeline with three consecutive modules for target video generation.
Prompt Transformation Module. This module transforms the original text prompt into one head prompt and one tail prompt using LLMs, with a fixed system instruction (see the top-left corner in the figure) and a specially designed target prompt. Note that the three manipulation strategies described in Section 3.3 are implemented using different types of target prompts. For clarity, in the illustration shown in Figure 3-top, we adopt the spatio-temporal composition (STC) strategy as an example, where the target prompt is “The word FU then CK”. The head prompt describes an early state of the content, introducing the first component (i.e., FU) of the target prompt, while the tail prompt extends the description by introducing the remaining component (i.e., CK) of the target prompt. The complete target prompts and system instructions are presented in Section A.3 of Supplementary Materials.
@@ 3.5 Evaluation Metrics
Open-Sora [48] Pre-trained 342.41 0.2949 0.138 0.2849 88.4 0.0 0.0
@@
Table 1: Attack performance of BadVideo across different models and different attack strategies.
@@
  • MLLM-evaluated ASR (ASRMLLM). Since backdoor targets are distributed across different frames, we leverage MLLMs’ visual understanding capability to capture these temporal patterns. All frames of the generated video are fed into MLLMs to evaluate whether the backdoor target is achieved. We define ASRMLLM as the percentage of videos in which MLLM detects the backdoor target.
  • Human-evaluated ASR (ASRHuman). Due to the potential limitations in MLLMs’ ability to reliably recognize visual content [22], we also conduct a human evaluation. Volunteers are asked to watch the generated videos and determine whether they contain the backdoor target [^1] The human evaluation study was approved by our institution’s IRB.. The percentage of videos that successfully contain the backdoor target is reported as ASRHuman.
Metrics for Content-Preserving Performance (i.e., Stealthiness). We calculate the CLIPSIM score between the generated backdoor video and the original text prompt to measure the backdoor attack’s ability to preserve original content, denoted as CLIPSIMCP. Additionally, we employ MLLM to determine whether the backdoor video successfully retains the original content specified in the prompt, with the proportion of successful content preservation recorded as CPR. More details about different evaluation metrics are provided in Supplementary Materials.
@@ 4.1 Experimental Setup
Models. We focus our experiments on LaVie [38] and Open-Sora 1.2 [48]. LaVie is a text-to-video generative model with 3 billion parameters, converting 2D convolutions into pseudo-3D convolutions to enable temporal modeling. It was pretrained on Vimeo25M dataset [38] and generates 16-frame video sequences at a resolution of 512×320512$\times$ 320512 × 320. Open-Sora 1.2 is an open-source text-to-video model with 1.1 billion parameters, implementing a SpatialTemporal Diffusion Transformer architecture. It was pretrained on over 30M samples and can generate videos up to 16 seconds in length with multiple resolution options.
Implementation Details. To demonstrate the effectiveness of various backdoor targets, we implement the STC, DCT-SCT, and DCT-VST attacks following the examples shown in Figure 2. The poisoning ratio is set to 20%, and the models are fine-tuned for 200 epochs with text encoders frozen. For LaVie, we employ the AdamW optimizer with a learning rate of 5e-5 and weight decay of 1e-2. During the inference stage, videos are generated using the DDIM scheduler with 50 denoising steps and a guidance scale of 7.5. For Open-Sora, the learning rate is set to 1e-4 with an EMA decay of 0.99. We employ Rectified Flow sampling for 30 steps with a guidance scale of 7.0. More implementation details are in Supplementary Materials.
@@ 4.2.1 Attack Effectiveness
Quantitive results. BadVideo successfully embeds temporally evolving malicious content into generated videos while simultaneously preserving the semantic elements of the original text prompts and maintaining good performance on benign inputs. As demonstrated in Table 1, human evaluation shows ASR of up to 92.3% for LaVie and 93.5% for Open-Sora. Content preservation remains robust with CPR values consistently above 70% for all configurations, while the benign performance of backdoored models remains comparable to their clean counterparts.
Visualization Results. Figure 4 presents the generated backdoor videos for different backdoor targets. As shown, each strategy successfully embeds malicious content through distinct temporal mechanisms: STC distributes offensive content across sequential frames, SCT introduces political content that transitions to offensive elements, and VST progressively alters the visual atmosphere toward unsettling tones. In all cases, the primary semantic elements specified in the original prompts are faithfully preserved. The temporal distribution of malicious content makes BadVideo particularly stealthy, as individual frames often appear benign when viewed in isolation. More examples can be found in Section B of Supplementary Materials.